# An Open-source Hardware/Software Architecture and Supporting Simulation Environment to Perform Human FPV Flight Demonstrations for Unmanned Aerial Vehicle Autonomy

### v1

@@ I INTRODUCTION
Unmanned Aerial Vehicle (UAV) refers to a variety of flying equipment capable of performing tasks without an onboard human pilot. Initially introduced in the early 19th century for military use, quadcopters have evolved significantly with the advancement of sensors and microprocessors [1]. Modern UAVs range from small, electrically powered devices suitable for indoor navigation to larger aircraft capable of carrying passengers on long-range missions. Among UAVs, quadcopters are particularly popular due to their versatility and stability. The transportation sector greatly benefits from quadcopters, especially in Urban Air Mobility (UAM), last-mile delivery, and traffic monitoring, where their abilities in accurate GPS navigation, obstacle avoidance, and payload handling are indispensable. In the realm of entertainment, quadcopters elevate aerial filming with multi-UAV cooperation[2], introducing both redundancy and challenges like increased collision risks and the need for advanced obstacle avoidance[3][4]. Their role in localization, particularly in search and rescue missions[5][6], is pivotal, addressing crises in diverse terrains and demanding exceptional maneuverability[7]. In construction, quadcopters are revolutionizing practices by enabling high-altitude visual inspections[8], pipeline monitoring[9], and site surveillance[10]. While the adoption of quadcopters for these complex and dangerous operational tasks significantly reduces the likelihood of human injury, it concurrently elevates the demand for highly skilled First Person View (FPV) quadcopter pilots who can manage these complex tasks. Market research projects a substantial growth in this demand, anticipating a 51.1 percent increase from 2022 to 2027[11]. In FPV flying, pilots control the quadcopters remotely from a first-person view using cameras mounted on them to navigate. In Urban Air Mobility (UAM) and last-mile delivery scenarios, FPV pilots expertly navigate through dense urban environments, where their ability to maneuver with precision is crucial for ensuring the safety of both people and the quadcopter themselves. The explosive advancement in artificial intelligence (AI) research further underscores this trend, shifting quadcopter operations towards increased autonomy and thereby diminishing the need for direct human pilot intervention [12]. This shift is particularly consequential for developing and operating electric Vertical Take-Off and Landing (VTOL) vehicles. AI’s capability to process vast amounts of data in real-time potentially enhances the safety and reliability of these vehicles. For VTOLs, which are complex systems requiring precise control, AI algorithms can significantly improve navigation accuracy, energy efficiency, and overall operational effectiveness. This evolution marks a pivotal transition from traditional piloted systems to more autonomous, AI-driven solutions, reshaping the future of urban air mobility and beyond. Current research largely focuses on the design and control of VTOL systems [13, 14, 15], operational efficiency, and safety measures, with few emerging studies examining the strategic use of reinforcement learning for air traffic control and scheduling within vertiports to optimize the deployment of eVTOLs in urban air mobility scenarios [16, 17, 18]. To effectively integrate autonomous capability into any aerial vehicle, including VTOL, its AI must be trained on extensive data derived from human pilots. This is critical for replicating human-like decision-making and providing a benchmark for autonomous flight systems. This ensures AI’s adaptability and public trust [19, 20]. Recent fields of research to align the performance of autonomous quadcopters with that of expert FPV quadcopter pilots refer to deep reinforcement learning [21]and imitation learning [22]. The importance of deep reinforcement learning in quadcopter autonomy is crucial, particularly for autonomous decision-making in complex environments. This is evident in research by Wang et al.[23], Çetin et al.[24], and Anwar and Raychowdhury[25], which demonstrates the effectiveness of reinforcement learning in improving quadcopters’ decision-making, especially in challenging settings like obstacle-rich areas or resource-limited environments. In the field of imitation learning, quadcopter learning from expert pilots significantly enhances maneuverability and efficiency. Most of these methods used existing open-source flight data or expensive hardware to collect new data, limiting access to those with substantial resources. This barrier restricts smaller research teams and individual innovators who may not have the financial capability to invest in high-cost equipment. Further, current open-source datasets like EuRoC [26] and Zurich Urban MAV [27] provide robust image-based data with flight state information, but they are limited by their third-person view collection methodology. Moreover, frameworks used for autonomous FPV racing rely on indoor motion capture systems, and outdoor versions require external Leica trackers [28] for ground truth location, significantly posing challenges in terms of cost and spatial limitations. Hence, the primary objective of this paper is to introduce an open-source framework equipped with FPV capabilities that is both cost-effective and readily accessible. This framework’s software stack aims to enable precise capturing of human pilot data, facilitating the development of advanced AI systems. The platform is based entirely on the open-source flight controller - Pixhawk. Pixhawk supports multiple flying modes, accommodating FPV Quadcopter pilots of varying experience levels with its dual-view FPV user interface. In conjunction with this interface, Mavlink [29] communication will be established to fetch real-time quadcopter state and control signal data during FPV flights. Additionally, this data will be synchronously saved with dual-frame images and corresponding timestamps after each flight. The advantage of using Pixhawk as a flight controller is its flexibility in switching between quadcopter frames, allowing an extendable change from quadcopter FPV to fixed-wing FPV quadcopter at an acceptable cost compared to current market camera quadcopter models. The secondary objective is to develop a digital twin platform (virtual environment) with Hardware-in-the-loop (HITL) capabilities and a GUI, providing an alternative solution for generating realistic, high-fidelity data, especially in scenarios where physical experimentation is expensive or challenging to perform very often [30]. This digital twin approach allows for safe, repeatable, and controlled testing without the constraints associated with physical experiments.
@@
Component Name Count Weight (g) Dimensions ($m$)
@@ II Hardware and Simulation Framework
The in-house built FPV quadcopter framework, as shown in figure 1, requires a finely-tuned FPV platform and an informative Graphical User Interface (GUI) to support pilots by providing real-time FPV camera views and flying state data such as ground speed, climbing speed, and relative altitude. We also implemented Airsim HITL with GUI in Unreal Engine to support FPV flight in virtual environment. The below subsections will deep into the development of hardware and software platforms. For simplicity, in later sections, the in-house built FPV quadcopter will be referred to as the physical quadcopter, and the Airsim HITL platform with compiled GUI will be referred to as the digital twin
@@ II-A4 Flight controller and companion computer
The flight controller is a critical component within this system, particularly given the aim of this study to develop a methodology for bridging the gap between the physical quadcopter framework and the digital twin. The Unreal Engine has achieved high-fidelity quadcopter simulation using AirSim’s HITL capability. Therefore, a flight controller that supports both HITL configurations and actual flight operations is essential. The Pixhawk 2.4.8 has been selected as the flight controller for the FPV quadcopter due to its affordability and capability to provide smooth flight control. Given the limitations of telemetry in the transmission rate of flight data, a Raspberry Pi has been integrated as a companion computer with the Pixhawk via a serial connection. This setup facilitates further development of the FPV GUI and enhances data collection capabilities
@@ II-C AirSim Simulation Environment
II-C1 Introduction to Simulation Environment
Microsoft AirSim [34] is an open-source robotics simulation platform. AirSim helps us solve the need for large data sets for training and allows debugging in a simulator. AirSim leverages current game engine rendering, physics, and perception computation to create accurate, real-world simulations. Together, this realism, based on efficiently generated ground-truth data, enables the study and execution of complex, time-consuming, and risky missions in the real world. The default quadcopter model in AirSim is modified to match our real hardware; the morphology and the cameras are modified to match the physical quadcopter. Airsim allows HITL simulation with PX4, an open-source, popular flight controller. HITL simulation offers a blend of simulation safety and real-world unpredictability. While software simulations are insightful, HITL simulation captures the nuances of actual hardware. At the same time, the HITL simulation starts, a python file will run to help retrieve the digital twin’s flying states along with image frames from two views, front and bottom, in corresponding timestamps simultaneously through Mavlink [29] and Airsim API [34]. A virtual version of the 24,000-square-foot Structure of Outdoor Autonomy Research (SOAR) facility at the University at Buffalo has been implemented through Unreal Engine 4.27. The environment is built precisely to match the real-world dimensions. In this virtual environment, pilots can free roam around using a graphic user interface (GUI) which includes two camera feeds, front view and bottom view, altitude information, and a top view global map, or there are few pre-defined tasks that the researchers can use while collecting data. Unreal Engine allows easy drag-and-drop features for creating a new task, which helps generate various tasks quickly to generate data.

### v2

@@ I INTRODUCTION
Unmanned Aerial Vehicle (UAV) refers to a variety of flying equipment capable of performing tasks without an onboard human pilot. Initially introduced in the early 19th century for military use, quadcopters have evolved significantly with the advancement of sensors and microprocessors [1]. Modern UAVs range from small, electrically powered devices suitable for indoor navigation to larger aircraft capable of carrying passengers on long-range missions. Among UAVs, quadcopters are particularly popular due to their versatility and stability. The transportation sector greatly benefits from quadcopters, especially in Urban Air Mobility (UAM), last-mile delivery, and traffic monitoring, where their abilities in accurate GPS navigation, obstacle avoidance, and payload handling are indispensable. In the realm of entertainment, quadcopters elevate aerial filming with multi-UAV cooperation [2], introducing both redundancy and challenges like increased collision risks and the need for advanced obstacle avoidance [3][4]. Their role in localization, particularly in search and rescue missions [5][6], is pivotal, addressing crises in diverse terrains and demanding exceptional maneuverability [7]. In construction, quadcopters are revolutionizing practices by enabling high-altitude visual inspections [8], pipeline monitoring [9], and site surveillance [10]. While the adoption of quadcopters for these complex and dangerous operational tasks significantly reduces the likelihood of human injury, it concurrently elevates the demand for highly skilled First Person View (FPV) quadcopter pilots who can manage these complex tasks. Market research projects a substantial growth in this demand, anticipating a 51.1 percent increase from 2022 to 2027 [11]. In FPV flying, pilots control the quadcopters remotely from a first-person view using cameras mounted on them to navigate. In Urban Air Mobility (UAM) and last-mile delivery scenarios, FPV pilots expertly navigate through dense urban environments, where their ability to maneuver with precision is crucial for ensuring the safety of both people and the quadcopter themselves.
@@
The explosive advancement in artificial intelligence (AI) research further underscores this trend, shifting quadcopter operations towards increased autonomy and thereby diminishing the need for direct human pilot intervention [12]. This shift is particularly consequential for developing and operating electric Vertical Take-Off and Landing (VTOL) vehicles. AI’s capability to process vast amounts of data in real-time potentially enhances the safety and reliability of these vehicles. For VTOLs, which are complex systems requiring precise control, AI algorithms can significantly improve navigation accuracy, energy efficiency, and overall operational effectiveness. This evolution marks a pivotal transition from traditional piloted systems to more autonomous, AI-driven solutions, reshaping the future of urban air mobility and beyond. Current research largely focuses on the design and control of VTOL systems [13, 14, 15], operational efficiency, and safety measures, with few emerging studies examining the strategic use of reinforcement learning for air traffic control and scheduling within vertiports to optimize the deployment of eVTOLs in urban air mobility scenarios [16, 17, 18]. To effectively integrate autonomous capability into any aerial vehicle, including VTOL, its AI must be trained on extensive data derived from human pilots. This is critical for replicating human-like decision-making and providing a benchmark for autonomous flight systems. This ensures AI’s adaptability and public trust [19, 20]. Recent fields of research to align the performance of autonomous quadcopters with that of expert FPV quadcopter pilots refer to deep reinforcement learning [21] and imitation learning [22]. The importance of deep reinforcement learning in quadcopter autonomy is crucial, particularly for autonomous decision-making in complex environments. This is evident in research by Wang et al. [23], Çetin et al. [24], and Anwar and Raychowdhury [25], which demonstrates the effectiveness of reinforcement learning in improving quadcopters’ decision-making, especially in challenging settings like obstacle-rich areas or resource-limited environments. In the field of imitation learning, quadcopter learning from expert pilots significantly enhances maneuverability and efficiency. Most of these methods used existing open-source flight data or expensive hardware to collect new data, limiting access to those with substantial resources. This barrier restricts smaller research teams and individual innovators who may not have the financial capability to invest in high-cost equipment. Further, current open-source datasets like EuRoC [26] and Zurich Urban MAV [27] provide robust image-based data with flight state information, but they are limited by their third-person view collection methodology. Moreover, frameworks used for autonomous FPV racing rely on indoor motion capture systems, and outdoor versions require external Leica trackers [28] for ground truth location, significantly posing challenges in terms of cost and spatial limitations.
Hence, the primary objective of this paper is to introduce an open-source framework equipped with FPV capabilities that is both cost-effective and readily accessible. This framework’s software stack aims to enable precise capturing of human pilot data, facilitating the development of advanced AI systems. The platform is based entirely on the open-source flight controller - Pixhawk. Pixhawk supports multiple flying modes, accommodating FPV Quadcopter pilots of varying experience levels with its dual-view FPV user interface. In conjunction with this interface, Mavlink [29] communication will be established to fetch real-time quadcopter state and control signal data during FPV flights. Additionally, this data will be synchronously saved with dual-frame images and corresponding timestamps after each flight. The advantage of using Pixhawk as a flight controller is its flexibility in switching between quadcopter frames, allowing an extendable change from quadcopter FPV to fixed-wing FPV quadcopter at an acceptable cost compared to current market camera quadcopter models. The secondary objective is to develop a digital twin platform (virtual environment) with Hardware-in-the-loop (HITL) capabilities and a GUI, providing an alternative solution for generating realistic, high-fidelity data, especially in scenarios where physical experimentation is expensive or challenging to perform very often [30]. This digital twin approach allows for safe, repeatable, and controlled testing without the constraints associated with physical experiments.
@@
Component Name Count Weight (g) Dimensions (mm)
@@ II Hardware and Simulation Framework
The in-house built FPV quadcopter framework, as shown in figure 1, requires a finely-tuned FPV platform and an informative Graphical User Interface (GUI) to support pilots by providing real-time FPV camera views and flying state data such as ground speed, climbing speed, and relative altitude. We also implemented Airsim HITL with GUI in Unreal Engine to support FPV flight in virtual environment. The below subsections will deep into the development of hardware and software platforms. For simplicity, in later sections, the in-house built FPV quadcopter will be referred to as the physical quadcopter, and the Airsim HITL platform with compiled GUI will be referred to as the digital twin.
@@ II-A4 Flight controller and companion computer
The flight controller is a critical component within this system, particularly given the aim of this study to develop a methodology for bridging the gap between the physical quadcopter framework and the digital twin. The Unreal Engine has achieved high-fidelity quadcopter simulation using AirSim’s HITL capability. Therefore, a flight controller that supports both HITL configurations and actual flight operations is essential. The Pixhawk 2.4.8 has been selected as the flight controller for the FPV quadcopter due to its affordability and capability to provide smooth flight control. Given the limitations of telemetry in the transmission rate of flight data, a Raspberry Pi has been integrated as a companion computer with the Pixhawk via a serial connection. This setup facilitates further development of the FPV GUI and enhances data collection capabilities.
@@ II-C AirSim Simulation Environment
II-C1 Introduction to Simulation Environment
Microsoft AirSim [34] is an open-source robotics simulation platform. AirSim helps us solve the need for large data sets for training and allows debugging in a simulator. AirSim leverages current game engine rendering, physics, and perception computation to create accurate, real-world simulations. Together, this realism, based on efficiently generated ground-truth data, enables the study and execution of complex, time-consuming, and risky missions in the real world. The default quadcopter model in AirSim is modified to match our real hardware; the morphology and the cameras are modified to match the physical quadcopter. Airsim allows HITL simulation with PX4, an open-source, popular flight controller. HITL simulation offers a blend of simulation safety and real-world unpredictability. While software simulations are insightful, HITL simulation captures the nuances of actual hardware. At the same time, the HITL simulation starts, a python file will run to help retrieve the digital twin’s flying states along with image frames from two views, front and bottom, in corresponding timestamps simultaneously through Mavlink [29] and Airsim API [34].
A virtual version of the 24,000-square-foot Structure of Outdoor Autonomy Research (SOAR) facility at the University at Buffalo has been implemented through Unreal Engine 4.27. The environment is built precisely to match the real-world dimensions. In this virtual environment, pilots can free roam around using a graphic user interface (GUI) which includes two camera feeds, front view and bottom view, altitude information, and a top view global map, or there are few pre-defined tasks that the researchers can use while collecting data. Unreal Engine allows easy drag-and-drop features for creating a new task, which helps generate various tasks quickly to generate data.
@@ V CONCLUSIONS
This platform can be useful to rigorously test and validate new AI-based algorithms in a safe and realistic simulation environment before deploying them on physical quadcopters or larger aircraft, which can reduce the risk of crashing in real-world testing and shorten the development time. We have also proposed a custom environment as an example for implementing collected data in RL algorithms. To improve the quality of data fed to the RL environment, future work could involve exploring methods to enhance the frequency of data acquisition and adding additional modules to verify the accuracy of the saved GPS coordinates. Thus, the demonstrated capabilities and initial flight test findings thereof points to both the need for and the feasibility of building open-source hardware/software platforms and digital twins to reduce barriers to entry for (and push forward) research in aerial autonomy.